Mudvayne, Inc., is trying to determine its cost of debt. The firm has a debt issue outstanding with 10 years to maturity that is quoted at 110 percent of face value. The issue makes semiannual payments and has an embedded cost of 8 percent annually.

What is the company’s pretax cost of debt? (Do not round intermediate calculation and round your answer to 2 decimal places. (e.g., 32.16))

  Cost of debt %

If the tax rate is 35 percent, what is the aftertax cost of debt? (Do not round intermediate calculations and round your answer to 2 decimal places. (e.g., 32.16))

  Cost of debt %
Answer #1

1)

Assuming face value to be $1,000

Number of periods = 10 * 2 = 20

Coupon = (0.08 * 1000) / 2 = 40

Price = 110% of 1,000 = 1,100

Pre tax cost of debt = 6.62%

Keys to use in a financial calculator: 2nd I/Y, FV 1000, PV -1100, PMT 40, N 20, CPT I/Y

2)

After tax cost if debt = 0.0662 (1 - 0.35)

After tax cost if debt = 0.0430 or 4.30%